# Massless higher-spin supermultiplets in 5D harmonic superspace

### v1

@@ 3.1 Prepotentials
5D ${\cal N}=1$ harmonic superspace with central charge may be parametrized by the coordinates $(x^{m},\theta^{\pm}_{\hat{\alpha}},u^{\pm}_{i},x^{5})$, where $m=0,1,2,3,4$ is the Lorentz index while $\hat{\alpha}=1,2,3,4$ is the spinor index; $x^{5}$ is the central charge coordinate. The $SU(2)_{R}$ harmonics $u^{\pm}_{i}$ obey the defining relations $u^{+i}u^{-}_{j}-u^{-i}u^{+}_{j}=\delta^{i}_{j}$, $u^{+i}u^{+}_{i}=u^{-i}u^{-}_{i}=0$. Associated with the harmonic coordinates are the covariant harmonic derivatives $^{++}, DD^{--}$ and $D^{0}$ with commutation relations of the $su(2)$ Lie algebra
@@
By analogy with the 4D case Galperin:1987em ; Galperin:1987ek , spin-2 supermultiplet in harmonic superspace may be described by a set of prepotentials $h^{++M}:=(h^{++m},h^{++\hat{\alpha}+},h^{++5})$ which comprise the following differential operator [^11] Associated with the covariant spinor derivative $D^{+}_{\hat{\alpha}}$ there is also a prepotential $h^{++\hat{\alpha}-}$. This prepotential, however, plays no role in the supergravity model because it can be safely gauged away. We will ignore analogous prepotentials in our construction of actions for massless higher spins in the 5D harmonic superspace.
@@
In the harmonic superspace formulation of supergravity, $h^{++M}$-prepotentials have a clear geometric meaning: they represent vielbeins for (super)diffeomorphism-covariant harmonic derivative,
@@
Covariant harmonic derivatives $\cal D$$^{++}$ and $\cal D$$^{--}$ should obey the commutation relations of $su(2)$ Lie algebra isomorphic to (9):
@@
For studies of linearized supergravity, it is sufficient to consider the zero-curvature condition (22) to the zeroth order in the coupling constant $\kappa$. In this case, this equation may written in the unfolded form
@@
It is straightforward to verify that these expressions solve for the equations (23) using the properties of the harmonic distributions described, e.g., in Refs. GIOS ; Galperin:1985bj . The solutions (24) represent spin-2 counterparts of analogous solutions to the harmonic zero-curvature condition in Yang-Mills theory in harmonic superspace Z ; Z2 . It is also possible to show that the gauge transformations (14) imply the following transformation laws for the prepotentials (24):
@@ 3.2 Unconstrained prepotentials
Supergravity prepotentials $h^{++M}$ introduced above obey the constraints (12). These constraint may be solved in terms of two unconstrained prepotentials, a full-superspace superfield $\Psi^{\hat{\alpha}-}$, and an analytic superfield $v^{++5}$, $D^{+}_{\hat{\alpha}}v^{++5}=0$:
@@
with unconstrained general superfield $l^{--\hat{\alpha}-}$ and an analytic superfield $l^{5}$, $D^{+}_{\hat{\alpha}}l^{5}=0$. In terms of these superfields, the $\lambda$-gauge transformations of the prepotentials $\Psi^{\hat{\alpha}-}$ and $v^{++5}$ are
@@
These variations imply gauge transformations (14).
As will be shown below, the unconstrained prepotentials will be necessary for deriving superfield equations of motions in the linearized supergravity and for obtaining conserved currents for $q$-hypermultiplet. The gauge symmetries (27) and (29) will play crucial role in obtaining corresponding conservation equations.
3.3 Action for linearized Poincaré supergravity
@@
Here $\mathrm{d}^{5|8}z\mathrm{d}u$ is the full superspace measure which is related with the analytic superspace measure $\mathrm{d}\zeta^{(-4)}$ in the standard manner,
@@
This is a linearized version of the Poincaré supergravity action (4). It would be interesting to develop a fully non-linear generalization of the supergravity action in the harmonic superspace which would reduce to (36).
@@ 3.4 Equations of motion
Although the general variation of the action (36) has simple form,
@@
derivation of the corresponding equations of motion is a non-trivial problem because the prepotentials $h^{++M}$ obey the constraints (12). The general solution of these constraints in terms of unconstrained superfields $\Psi^{\hat{\alpha}-}$ and $v^{++5}$ is given in Eqs. (26). Substituting Eqs. (26) into Eq. (37) we find the variation of the action (36) in terms of the unconstrained prepotentials:
@@
As a result, the equations of motion are
$0=$ $\frac{\delta S}{\delta\Psi^{\hat{\alpha}-}}=\frac{\mathrm{i}}{2}\Gamma^{m}_{\hat{\alpha}\hat{\beta}}(D^{+})^{2}D^{+\hat{\beta}}h^{--}_{m}-\frac{3\mathrm{i}}{2}(D^{+})^{2}D^{+}_{\hat{\alpha}}h^{--5}\,$, (39a)
$0=$ $\frac{\delta S}{\delta v^{++5}}=-6(D^{+})^{4}h^{--5}\,.$ (39b)
The latter equation follows from the former one upon application of the derivative $D^{+\hat{\alpha}}$. Using the relations (15) it is possible to show that the equations (39) remain invariant under $\lambda$-gauge transformations of prepotentials (25).
Equation (39a) may be represented in the form of analyticity condition for a scalar superfield $W$:
$D^{+}_{\hat{\alpha}}W=0\,$, (40)
@@
$W:=D^{+}_{\hat{\alpha}}D^{+}_{\hat{\beta}}\Gamma_{m}^{\hat{\alpha}\hat{\beta}}h^{--m}+3(D^{+})^{2}h^{--5}\,.$ (41)
This superfield gets shifted under $\lambda$-gauge transformation (25), $\delta_{\lambda}W=-8\mathrm{i}\Omega$, where $\Omega=\partial_{m}\lambda^{m}-D^{-}_{\hat{\alpha}}\lambda^{\hat{\alpha}+}$ is an analytic superfield.
@@ 3.5.1 Gauge-invariant cubic vertex
The power of this formulation is that it naturally extends to include the interaction with background spin-2 supermultiplet: it is sufficient to promote the harmonic derivative $D^{++}$ to the covariant derivative $\cal D$$^{++}$ given by Eq. (18). As a result, the cubic interaction vertex of the hypermultiplet with the background spin-2 supermultiplet reads
@@
This variation is a non-linear extension of the gauge transformation (14). In what follows we will restrict ourselves to the linear gauge transformations of the prepotentials, as the construction of fully non-linear supergravity model leads us too far from the goals of this paper.
@@ 3.5.2 Equations of motion and conserved currents
Unlike prepotentials $h^{++M}$, the hypermultiplet may have non-trivial dependence on the central charge coordinate $x^{5}$ if it has a non-vanishing mass $m$. This dependence, however, may be only cyclic,
@@
Free hypermultiplet equation of motion
@@
implies $D^{--}D^{--}q^{+a}=0$. Acting by the derivatives $D^{+}_{\hat{\alpha}}$ on the latter equation it is possible to extract its differential consequences:
@@
where $\Gamma^{++}$ is connection (which should not be confused with the $SO(4,1)$ $\Gamma$-matrices $\Gamma^{m}$)
@@
Under gauge transformations (14) this superfield varies as
@@
These currents are conserved on free hypermultiplet equation of motion:
@@
These conservation conditions follow from the $\lambda$-gauge symmetry of the prepotentials (29). $\rho$-gauge symmetry (27) in its turn dictates the following conservation laws for the currents:
@@
In the case of the free hypermultiplet model, the scalar supercurrent has simple form
@@
In 4D field theories, the scalar supercurrent $\cal J$ and the multiplet of anomalies $\cal T$$^{++}$ were found in Ref. KT ; Butter:2010sc . These superfields are harmonic superspace counterparts of the supercurrent constructed by Sohnius Sohnius-supercurrent in conventional superspace approach. In Appendix C.4 we show that the supercurrent (64) originates from the $q$-hypermultiplet model coupled to the conformal supergravity background.
@@ 3.6 Component structure of prepotentials
The component structure of the prepotentials $h^{++M}$ may be suitably studied in the analytic basis with coordinates $(x^{m}_{A},\theta^{+}_{\hat{\alpha}},u^{\pm}_{i},x^{5}_{A})$ which are related with the central basis $(x^{m},\theta^{\pm}_{\hat{\alpha}},u^{\pm}_{i},x^{5})$ through the identities
@@
In this basis, the analyticity constraint is satisfied automatically because the corresponding covariant spinor derivative is short, $D^{+}_{\hat{\alpha}}=\frac{\partial}{\partial\theta^{-\hat{\alpha}}}$. Explicit expressions for other spinor and harmonic derivatives are given by Eqs. (196) and (197).
@@ 3.6.1 Bosonic fields
By analyzing these relations it is easy to see that the gauge freedom in $l_{1}$, $l_{2}^{m}$ and $l_{3}^{mn}$ may be used to gauge away some of the remaining auxiliary fields and achieve standard gauge variations for the physical ones:
@@
As a result, we conclude that, modulo gauge freedom associated with $b_{m}$ and $a$, the physical components $B_{(mn)}$ and $A_{m}$ as well as the auxiliary ones $V^{m\,ij}$, $D^{ij}$ and $P^{\hat{\alpha}\hat{\beta}}$ comprise 48 bosonic degrees of freedom (d.o.f.) off shell. Quadratic action for these fields is given in Appendix B.3, see Eq. (318).
@@ 3.6.2 Fermionic fields
Substituting these decompositions into Eqs. (70) we find residual gauge transformation of the fermionic components:
@@ 4.1 Prepotentials and gauge transformations
and obey the following algebraic constraints (Ansatz):
@@
    $h^{++Mm_{1}\ldots m_{s-2}}=h^{++(Mm_{1}\ldots m_{s-2})}\,;$ (89)
    (Here the symmetrization involves only the Lorentz index $m$ from the multi-index $M=(m,\hat{\alpha}+,5)$.)
@@
These constraints eliminate reducible (lower-spin) components from this multiplet and retain the irreducible spin-$s$ massless supermultiplet. As will be shown below, these constraints are consistent with gauge invariance and supersymmetry. Moreover, the constraint (91) will be crucial in constructing gauge-invariant quadratic action for higher spin supermultiplets.
@@
$\lambda^{Mm_{1}\ldots m_{s-2}}=\lambda^{(Mm_{1}\ldots m_{s-2})}\,,\qquad\eta_{m_{1}m_{2}}\lambda^{Mm_{1}m_{2}\ldots m_{s-2}}=0\,$, (97)
@@
Moreover, since the gauge transformation (95) should preserve the analyticity properties of prepotentials (93), the gauge parameters must obey analogous constraints:

### v2

@@ 3.1 Prepotentials
5D ${\cal N}=1$ harmonic superspace with central charge may be parametrized by the coordinates $(x^{m},\theta^{\pm}_{\hat{\alpha}},u^{\pm}_{i},x^{5})$, where $m=0,1,2,3,4$ is the Lorentz index while $\hat{\alpha}=1,2,3,4$ is the spinor index, and $x^{5}$ is the central charge coordinate. The $SU(2)_{R}$ harmonics $u^{\pm}_{i}$ obey the defining relations $u^{+i}u^{-}_{j}-u^{-i}u^{+}_{j}=\delta^{i}_{j}$, $u^{+i}u^{+}_{i}=u^{-i}u^{-}_{i}=0$. Associated with the harmonic coordinates are the covariant harmonic derivatives $^{++}, DD^{--}$ and $D^{0}$ with commutation relations of the $su(2)$ Lie algebra
@@
By analogy with the 4D case Galperin:1987em ; Galperin:1987ek , the spin-2 supermultiplet in harmonic superspace may be described by a set of prepotentials $h^{++M}:=(h^{++m},h^{++\hat{\alpha}+},h^{++5})$ which comprise the following differential operator [^14] Associated with the covariant spinor derivative $D^{+}_{\hat{\alpha}}$ there is also a prepotential $h^{++\hat{\alpha}-}$. This prepotential, however, plays no role in the supergravity model because it can be safely gauged away. We will ignore analogous prepotentials in our construction of actions for massless higher spins in the 5D harmonic superspace.
@@
Within the harmonic superspace formulation for supergravity, the prepotentials $h^{++M}$ have a clear geometric meaning: they represent vielbeins for a (super)diffeomorphism-covariant harmonic derivative,
@@
Covariant harmonic derivatives $\cal D$$^{++}$ and $\cal D$$^{--}$ should obey the commutation relations of the $su(2)$ Lie algebra isomorphic to (10):
@@
For studies of linearized supergravity, it is sufficient to consider the zero-curvature condition (23) to zeroth order in the coupling constant $\kappa$. In this case, this equation may written in the unfolded form
@@
It is straightforward to verify that these expressions solve the equations (24) by using the properties of the harmonic distributions described, e.g., in Refs. GIOS ; Galperin:1985bj . The solutions (25) represent spin-2 counterparts of analogous solutions to the harmonic zero-curvature condition in the Yang-Mills theory in harmonic superspace Z ; Z2 . It is also possible to show that the gauge transformations (15) imply the following transformation laws for the prepotentials (25):
@@ 3.2 Unconstrained prepotentials
The supergravity prepotentials $h^{++M}$ introduced above obey the constraints (13). These constraint may be solved in terms of two unconstrained prepotentials, a full-superspace superfield $\Psi^{\hat{\alpha}-}$, and an analytic superfield $v^{++5}$, $D^{+}_{\hat{\alpha}}v^{++5}=0$:
@@
with an unconstrained general superfield $l^{--\hat{\alpha}-}$ and an analytic superfield $l^{5}$, $D^{+}_{\hat{\alpha}}l^{5}=0$. In terms of these superfields, the $\lambda$-gauge transformations of the prepotentials $\Psi^{\hat{\alpha}-}$ and $v^{++5}$ are
@@
These variations imply the gauge transformations (15).
As will be shown below, the unconstrained prepotentials will be necessary for deriving superfield equations of motions in the linearized supergravity and for obtaining conserved currents for the $q$-hypermultiplet. The gauge symmetries (28) and (30) will play crucial role in obtaining corresponding conservation equations.
3.3 Action for the linearized Poincaré supergravity
@@
Here $\mathrm{d}^{5|8}z\mathrm{d}u$ is the full superspace measure which is related to the analytic superspace measure $\mathrm{d}\zeta^{(-4)}$ in the standard manner,
@@
This is a linearized version of the Poincaré supergravity action (4). It would be interesting to develop a non-linear generalization of the supergravity action in the harmonic superspace which reduces to (37).
@@ 3.4 Equations of motion
Although the general variation of the action (37) has a simple form,
@@
it is not completely trivial to derive the corresponding equations of motion, since the prepotentials $h^{++M}$ obey the constraints (13). However, the general solution of these constraints in terms of unconstrained superfields $\Psi^{\hat{\alpha}-}$ and $v^{++5}$ is given in Eqs. (27). Substituting Eqs. (27) into Eq. (38) we find the variation of the action (37) in terms of the unconstrained prepotentials:
@@
Let us consider the variational derivative of the action with respect to $\Psi^{\hat{\alpha}-}$,
$\Xi^{+}_{\hat{\alpha}}:=2\mathrm{i}\frac{\delta S}{\delta\Psi^{\hat{\alpha}-}}=(\Gamma^{m})_{\hat{\alpha}}{}^{\hat{\beta}}(D^{+})^{2}D^{+}_{\hat{\beta}}h^{--}_{m}+3(D^{+})^{2}D^{+}_{\hat{\alpha}}h^{--5}\,.$ (40)
This superfield proves to be gauge-invariant, $\delta_{\lambda}\Xi^{+}_{\hat{\alpha}}=0$, and obeys the equation
$D^{++}\Xi^{+}_{\hat{\alpha}}=0\quad\implies\quad\Xi^{+}_{\hat{\alpha}}(z,u)=\Xi^{i}_{\hat{\alpha}}(z)u_{i}^{+}~.$ (41)
The harmonic-independent superfield $\Xi^{i}_{\hat{\alpha}}$ is the linearized version of the curvature spinor introduced in section 2.
The equation of motion for $\Psi^{\hat{\alpha}-}$, which follows from (39), is
$\Xi^{+}_{\hat{\alpha}}=0\,.$ (42)
This is the linearized version of (5). It should be pointed out that $\Xi^{+}_{\hat{\alpha}}$ may be expressed in the form
$\Xi^{+}_{\hat{\alpha}}=D^{+}_{\hat{\alpha}}T\,$, (43)
@@
$T:=D^{+}_{\hat{\alpha}}D^{+}_{\hat{\beta}}\Gamma_{m}^{\hat{\alpha}\hat{\beta}}h^{--m}+3(D^{+})^{2}h^{--5}\,.$ (44)
This superfield gets shifted under the $\lambda$-gauge transformation (26), $\delta_{\lambda}T=-8\mathrm{i}\Omega$, where $\Omega=\partial_{m}\lambda^{m}-D^{-}_{\hat{\alpha}}\lambda^{\hat{\alpha}+}$ is an analytic superfield.
The equation of motion for $v^{++5}$, which follows from (39), is
$0=\frac{\delta S}{\delta v^{++5}}=-6(D^{+})^{4}h^{--5}\,.$ (45)
This equation is, however, not independent. It is obtained from Eq. (42) by applying the derivative $D^{+\hat{\alpha}}$.
3.5 Linearized super-Weyl tensor
The superfield description of the Weyl multiplet for conformal supergravity was developed in Ref. KT-M08 ; Butter:2014xxa . The super Weyl tensor is a primary dimension-1 antisymmetric tensor superfield, $W_{ab}=-W_{ba}$. In the linearized case, there is a unique (modulo an overall coefficient) gauge-invariant expression for this tensor in terms of the prepotentials (25):
$W_{ab}=$ $\,(D^{+})^{2}\partial_{[a}h_{b]}^{--}+\frac{1}{4}\varepsilon_{abcde}(\Gamma^{e})^{\hat{\alpha}\hat{\beta}}D^{+}_{\hat{\alpha}}D^{+}_{\hat{\beta}}\partial^{c}h^{--d}$
[…]
Using the identities (16) and (192) it is possible to prove that (46) is invariant under the $\lambda$-gauge transformations (26), $\delta_{\lambda}W_{ab}=0$. In a similar way, the identities (13) and (24) allow one to verify that this tensor is harmonic-independent, $D^{++}W_{ab}=0$.
The expression (46) is a 5D counterpart of the 4D ${\cal N}=2$ super-Weyl tensor constructed in harmonic superspace in Ref. Ivanov:2024gjo .
@@ 3.6.1 Gauge-invariant cubic vertex
The power of this formulation is that it naturally extends to include the interaction with the background spin-2 supermultiplet: it is sufficient to promote the harmonic derivative $D^{++}$ to the covariant derivative $\cal D$$^{++}$ given by Eq. (19). As a result, the cubic interaction vertex of the hypermultiplet with the background spin-2 supermultiplet reads
@@
This variation is a nonlinear extension of the gauge transformation (15). In what follows we will restrict ourselves to the linear gauge transformations of the prepotentials, as the construction of non-linear supergravity models leads us too far from the goals of this paper.
@@ 3.6.2 Equations of motion and conserved currents
Unlike the prepotentials $h^{++M}$, the hypermultiplet may have non-trivial dependence on the central charge coordinate $x^{5}$ if it has a non-vanishing mass $m$. This dependence, however, may only be cyclic,
@@
The free hypermultiplet equation of motion
@@
implies $D^{--}D^{--}q^{+a}=0$. By acting with the derivatives $D^{+}_{\hat{\alpha}}$ on the latter equation it is possible to extract its differential consequences:
@@
where $\Gamma^{++}$ is the connection (which should not be confused with the $SO(4,1)$ $\Gamma$-matrices $\Gamma^{m}$),
@@
Under the gauge transformations (15) this superfield varies as
@@
These currents are conserved on the free hypermultiplet equation of motion:
@@
These conservation conditions follow from the $\lambda$-gauge symmetry of the prepotentials (30). $\rho$-gauge symmetry (28) dictates the following conservation laws for the currents:
@@
In the case of the free hypermultiplet model, the scalar supercurrent has the simple form
@@
In 4D field theories, the scalar supercurrent $\cal J$ and the multiplet of anomalies $\cal T$$^{++}$ were found in Ref. KT ; Butter:2010sc . These superfields are related to the supercurrent constructed by Sohnius Sohnius-supercurrent . In Appendix C.4 we show that the supercurrent (69) originates from the $q$-hypermultiplet model coupled to the conformal supergravity background.
@@ 3.7 Component structure of prepotentials
The component structure of the prepotentials $h^{++M}$ can be studied in the analytic basis with coordinates $(x^{m}_{A},\theta^{+}_{\hat{\alpha}},u^{\pm}_{i},x^{5}_{A})$, which are related to the central basis $(x^{m},\theta^{\pm}_{\hat{\alpha}},u^{\pm}_{i},x^{5})$ through the identities
@@
In this basis, the analyticity constraint is satisfied because the corresponding covariant spinor derivative is short, $D^{+}_{\hat{\alpha}}=\frac{\partial}{\partial\theta^{-\hat{\alpha}}}$. Explicit expressions for the other spinor and harmonic derivatives are given by Eqs. (201) and (202).
@@ 3.7.1 Bosonic fields
By analyzing these relations it is easy to see that the gauge freedom in $l_{1}$, $l_{2}^{m}$ and $l_{3}^{mn}$ may be used to gauge away some of the remaining auxiliary fields and achieve the standard gauge variations for the physical ones:
@@
As a result, we conclude that, modulo the gauge freedom associated with $b_{m}$ and $a$, the physical components $B_{(mn)}$ and $A_{m}$ as well as the auxiliary ones $V^{m\,ij}$, $D^{ij}$ and $P^{\hat{\alpha}\hat{\beta}}$ comprise 48 bosonic degrees of freedom (d.o.f.) off shell. The quadratic action for these fields is given in Appendix B.3, see Eq. (323).
@@ 3.7.2 Fermionic fields
Substituting these decompositions into Eqs. (75), we find the residual gauge transformations of the fermionic components:
@@ 4.1 Prepotentials and gauge transformations
and are required to obey the following algebraic constraints:
@@
    $h^{++m_{1}\ldots m_{s-1}}=$ $\,h^{++(m_{1}\ldots m_{s-1})}\,$, (94a)
    $h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}}=$ $\,h^{++\hat{\alpha}+(m_{1}\ldots m_{s-2})}\,$, (94b)
    $h^{++5m_{1}\ldots m_{s-2}}=$ $\,h^{++5(m_{1}\ldots m_{s-2})}\,;$ (94c)
@@
These constraints eliminate reducible (lower-spin) components from this multiplet and retain the irreducible spin-$s$ massless supermultiplet. As will be shown below, these constraints are consistent with gauge invariance and supersymmetry. Moreover, the constraint (96) will be crucial in constructing a gauge-invariant quadratic action for higher spin supermultiplets.
@@
$\lambda^{m_{1}\ldots m_{s-1}}=$ $\,\lambda^{(m_{1}\ldots m_{s-1})}\,$, (102a)
$\lambda^{\hat{\alpha}+m_{1}\ldots m_{s-2}}=$ $\,\lambda^{\hat{\alpha}+(m_{1}\ldots m_{s-2})}\,$, (102b)
$\lambda^{5m_{1}\ldots m_{s-2}}=$ $\,\lambda^{5(m_{1}\ldots m_{s-2})}\,$, (102c)
$\eta_{m_{1}m_{2}}\lambda^{Mm_{1}m_{2}\ldots m_{s-2}}=$  $\,0\,$, (102d)
@@
Moreover, since the gauge transformation (100) should preserve the analyticity properties of the prepotentials (98), the gauge parameters must obey analogous constraints: